# Changing Answer Order Can Decrease MMLU Accuracy

### v1

@@ 1 Introduction
$\dagger$ This work was done during their internships at FAIR, Meta.
One of the largest outstanding issues with interpreting the results of model evaluation pertains to the robustness of accuracy measurements. For example, NLP model accuracy has been shown to be fairly brittle. For example, accuracy can drop when researchers apply input alterations based on paraphrasing (Gan and Ng, 2019), word order changes (Gauthier and Levy, 2019; Ribeiro et al., 2020; Sinha et al., 2021a, 2022; Allen-Zhu and Li, 2023a, b; Berglund et al., 2023; Golovneva et al., 2024; Kitouni et al., 2024), or other minor, largely meaning-preserving input variations or perturbations (Belinkov and Bisk, 2018; Ebrahimi et al., 2018; Jiang et al., 2020; Gao et al., 2021; Li et al., 2021; Sinha et al., 2021b; Moradi and Samwald, 2021; Papakipos and Bitton, 2022; Qian et al., 2022; Goodarzi et al., 2023; Sinha et al., 2023). If many models fail to be robust on a benchmark, regardless of their initially measured accuracy, we may need to reconsider how we use it as the basis for a leaderboard that actually ranks models.
@@ 3 Results
4 Discussion & Conclusion
Related Work.
Like all evaluation datasets, validity is important. Several recent works have discussed MMLU’s validity (Gema et al., 2024; Zheng et al., 2023; Wang et al., 2024a, b). In particular, Wang et al. (2024b) found trivial and noisy questions in the dataset and proposed an update, MMLU-Pro, which aims to mitigate those issues. Concurrent work on model robustness to question-answering order (Zhou et al., 2024) applies a similar approach to ours that shuffles answer label content and also explores other possible modes of interrogating robustness. While they also find non-robustness to question variants, our work differs from theirs in that our metric can account for the multiplicity of potential orderings of answer labels; we also provide further analysis for each category in MMLU in Figure 3 in the appendix.
Conclusion.

### v2

@@ Abstract
$\dagger$ This work was done during their internships at FAIR, Meta.
@@ 1 Introduction
One of the largest outstanding issues with interpreting the results of model evaluation pertains to the robustness of accuracy measurements. For example, NLP model accuracy has been shown to be fairly brittle. For example, accuracy can drop when researchers apply input alterations based on paraphrasing (Gan and Ng, 2019), word order changes (Gauthier and Levy, 2019; Ribeiro et al., 2020; Sinha et al., 2021a, 2022; Allen-Zhu and Li, 2023a, b; Berglund et al., 2023; Golovneva et al., 2024; Kitouni et al., 2024; Sugawara et al., 2020), or other minor, largely meaning-preserving input variations or perturbations (Belinkov and Bisk, 2018; Ebrahimi et al., 2018; Jiang et al., 2020; Gao et al., 2021; Li et al., 2021; Sinha et al., 2021b; Moradi and Samwald, 2021; Papakipos and Bitton, 2022; Qian et al., 2022; Goodarzi et al., 2023; Sinha et al., 2023). If many models fail to be robust on a benchmark, regardless of their initially measured accuracy, we may need to reconsider how we use it as the basis for a leaderboard that actually ranks models.
@@ 3 Results
4 Related Works
LLMs can be Sensitive to Option Order and Label.
Recent works have also shown that the accuracy of models on multiple-choice question datasets can change significantly when the order of answer options is rearranged Robinson and Wingate (2023); Pezeshkpour and Hruschka (2024); Alzahrani et al. (2024); Wei et al. (2024); Xue et al. (2024); Zong et al. (2024). This suggests that models are sensitive to the order of answer options, which can impact their performance. Wang et al. (2024a) studies how changing the number of options in mutliple choice question datasets affect the model performance. They found that LLMs have overfitted to multiple choice question datasets with exactly four options.
Other studies have shown that models may exhibit prior biases towards specific option IDs (e.g., ‘A’) Wei et al. (2024); Zheng et al. (2023b); Reif and Schwartz (2024); Ross et al. (2024); Li and Gao (2024); Zheng et al. (2023a). Some works have also shown that models can perform surprisingly well above random chance even when question text is removed and only answer options are provided Balepur et al. (2024); Shah et al. (2020); Balepur and Rudinger (2024). Recent works have also shown that replacing the correct option with “None of the Above” leads to a drastic decline in performance across all models (Wang et al., 2024a; Xu et al., 2024). These findings suggest that models may be relying on artifacts or biases in the data rather than truly understanding the questions Röttger et al. (2024); Raj et al. (2023).
In a concurrent work, McIlroy-Young et al. (2024) proposed a solution fro mitigating the issue of order dependency in LLMs by modifying the self attention matrix of the input sequence. They set the attention scores between different options to be zero, effectively preventing the model from attending to the order of options.
In contrast to above works, our work focuses on category-wise differences in model performance and proposes a new metric that takes into account the variation in model performance across different answer orders. Our approach provides a more nuanced understanding of the impact of answer option ordering on model accuracy.
Evaluation Dataset Validity.
For all evaluation datasets, validity is important, and MMLU is no exception. Several recent works have discussed MMLU’s validity (Gema et al., 2024; Zheng et al., 2023a; Wang et al., 2024a, b). In particular, Wang et al. (2024b) found trivial and noisy questions in the dataset and proposed an update, MMLU-Pro, which aims to mitigate those issues. Concurrent work on model robustness to question-answering order (Zhou et al., 2024) applies a similar approach to ours that shuffles answer label content and also explores other possible modes of interrogating robustness. While they also find non-robustness to question variants, our work differs from theirs in that our metric can account for the multiplicity of potential orderings of answer labels; we provide further analysis for each category in MMLU in Figure 3 in the appendix.
5 Conclusion